# Edge Importance in Complex Networks

### v1

@@ 2.2 Methods for large networks
Recently, Kandolf et al. [16] derived a method for evaluating approximations of the Fréchet derivative of a matrix function by Krylov subspace methods. Applications of this technique to network analysis have recently been discussed by De la Cruz Cabrera et al. [4] and Schweitzer [27]. We first outline this method and subsequently discuss some alternatives.
@@
where $\Gamma$ is a curve in $\Omega$ that winds around the spectrum of $\mathbf{A}$ exactly once and $\mathrm{i}=\sqrt{-1}$. In this paper, we are primarily interested in the situation when $f(z)=\exp_{0}(z)$, but the techniques discussed apply to other analytic functions as well. Let $u,v\in\mathbb{R}^{n}$ be nonvanishing vectors. Kandolf et al. [16] show that the Fréchet derivative of $f$ at $\mathbf{A}$ in the direction $uv^{T}$ can be expressed as
@@
and determine an approximation of this expression by using Krylov subspace techniques to approximate the vectors
@@
where the superscript H denotes transposition and complex conjugation. Kandolf et al. [16] and Schweitzer [27] approximate the vectors $s(z)$ and $t(z)$ by a Krylov subspace technique based on the Arnoldi process. Application of $1\leq m\ll n$ steps of the Arnoldi process to $\mathbf{A}$ with initial vector $u$, and to $\mathbf{A}^{T}$ with initial vector $v$, generically, yields the Arnoldi decompositions
$\mathbf{A}\mathbf{V}_{m}=\mathbf{V}_{m}\mathbf{G}_{m}+{\tilde{v}}_{m+1}e_{m}^{T},\qquad\mathbf{A}^{T}\mathbf{W}_{m}=\mathbf{W}_{m}\mathbf{H}_{m}+{\tilde{w}}_{m+1}e_{m}^{T},$ (11)
@@
where the matrices $\mathbf{G}_{m},\mathbf{H}_{m}\in\mathbb{R}^{m\times m}$ are of upper Hessenberg form, the matrix $\mathbf{V}_{m}\in\mathbb{R}^{n\times m}$ has orthonormal columns with initial column $u/\|u\|_{2}$, the vector ${\tilde{v}}_{m+1}\in\mathbb{R}^{n}$ is orthogonal to the range of $\mathbf{V}_{m}$, the matrix $\mathbf{W}_{m}\in\mathbb{R}^{n\times m}$ has orthonormal columns with initial column $v/\|v\|_{2}$, and the vector ${\tilde{w}}_{m+1}\in\mathbb{R}^{n}$ is orthogonal to the range of $\mathbf{W}_{m}$; see, e.g., Saad [26, Chapter 6] for details on the Arnoldi process. Here we only note that the evaluation of the decompositions (11) requires the evaluation of $m$ matrix-vector products with the matrix $\mathbf{A}$ and $m$ matrix-vector products with the matrix $\mathbf{A}^{T}$. This is the dominating computational work when the matrix $\mathbf{A}$ is large and the number of Arnoldi steps is fairly small. We assume here that the Arnoldi processes do not break down when computing (11); in case of breakdown, the formulas (11) simplify.
Kandolf et al. [16] propose to use the Arnoldi approximation
$L_{f,{\rm Arnoldi}}(\mathbf{A},uv^{T})=\mathbf{V}_{m}\mathbf{X}_{m}\mathbf{W}_{m}^{T},$ (12)
of $L_{f}(\mathbf{A},uv^{T})$, where $\mathbf{X}_{m}$ is the upper right $m\times m$ submatrix of the $2m\times 2m$ matrix
$f\left(\left[\begin{array}[]{cc}\mathbf{G}_{m}&\|u\|_{2}\|v\|_{2}e_{1}e_{1}^{T}\\ \mathbf{0}&\mathbf{H}_{m}^{T}\end{array}\right]\right)=\left[\begin{array}[]{cc}f(\mathbf{G}_{m})&\mathbf{X}_{m}\\ \mathbf{0}&f(\mathbf{H}_{m}^{T})\end{array}\right]$ (13)
@@
with $e_{1}=[1,0,\ldots,0]^{T}$. Schweitzer [27] applies formulas (11), (12), and (13) with $\mathbf{A}$ replaced by $\mathbf{A}^{T}$ and $u=v=e$; cf. (6). Convergence results are provided by Kandolf et al. [16].
@@
An alternative approach to compute an approximation of the Fréchet derivative $L_{f}(\mathbf{A}^{T},ee^{T})$ is to apply finite-difference approximations analogously as (10). Application of $m$ steps of the Arnoldi process to the matrix $\mathbf{A}^{T}$ with initial vector $e$ gives the Arnoldi decomposition
$\mathbf{A}^{T}\mathbf{V}_{m}=\mathbf{V}_{m}\mathbf{H}_{m}+{\tilde{v}}_{m+1}e_{m}^{T},$
@@
where the columns of $\mathbf{V}_{m}\in\mathbb{R}^{n\times m}$ are orthonormal and span the Krylov subspace
$\mathbf{K}_{m}(\mathbf{A}^{T},e)={\rm span}\{e,\mathbf{A}^{T}e,(\mathbf{A}^{T})^{2}e,\ldots,(\mathbf{A}^{T})^{m-1}e\}$
@@
and the first column of $\mathbf{V}_{m}$ is $e/\sqrt{n}$. Moreover, the vector ${\tilde{v}}_{m+1}\in\mathbb{R}^{n}$ is orthogonal to $\mathbf{K}_{m}(\mathbf{A}^{T},e)$ and $\mathbf{H}_{m}\in\mathbb{R}^{m\times m}$ is an upper Hessenberg matrix. Then
$\mathbf{V}_{m}^{T}\mathbf{A}^{T}\mathbf{V}_{m}=\mathbf{H}_{m}.$
@@
$\mathbf{A}^{T}\approx\mathbf{V}_{m}\mathbf{H}_{m}\mathbf{V}_{m}^{T}$ (14)
@@
$f(\mathbf{A}^{T})\approx\mathbf{V}_{m}f(\mathbf{H}_{m})\mathbf{V}_{m}^{T}.$ (15)
The approximation (15) is quite accurate when $f(\mathbf{A}^{T})$ can be approximated well by a matrix of low rank. This is the case for many real-life undirected networks when $f(t)=\exp(t)$; see [14]. The adjacency matrix $\mathbf{A}$ is for many directed graphs that arise in real-life applications quite close to symmetric. This suggests that the approximation (15) is fairly accurate for these kinds of graphs. Illustrations that this, indeed, is the case can be found in Section 4.
@@
$\mathbf{A}^{T}+hee^{T}\approx\mathbf{V}_{m}\mathbf{H}_{m}\mathbf{V}_{m}^{T}+hee^{T}=\mathbf{V}_{m}(\mathbf{H}_{m}+hne_{1}e_{1}^{T})\mathbf{V}_{m}^{T}$
@@
$f(\mathbf{A}^{T}+hee^{T})\approx\mathbf{V}_{m}f(\mathbf{H}_{m}+hne_{1}e_{1}^{T})\mathbf{V}_{m}^{T},$
@@
$\frac{f(\mathbf{A}^{T}+hee^{T})-f(\mathbf{A}^{T}-hee^{T})}{2h}\approx\mathbf{V}_{m}\frac{f(\mathbf{H}_{m}+hne_{1}e_{1}^{T})-f(\mathbf{H}_{m}-hne_{1}e_{1}^{T})}{2h}\mathbf{V}_{m}^{T}.$ (16)
We will use the expression on the right-hand side as an approximation of $L_{f}(\mathbf{A}^{T},ee^{T})$ in computed examples with $h$ the same as in Example 2.1. Note that the evaluation of this expression only requires the computation of $m$ matrix-vector products with the matrix $\mathbf{A}^{T}$. For large-scale problems for which the evaluation of matrix-vector products is the dominant computational work, the use of the right-hand side of (16) halves the computational burden when compared with the evaluation of (12).
We turn to the situation when the matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$ is symmetric and first review the computations described by Kandolf et al. [16] of the analogue of the expression (12) when the direction is $ee^{T}$. Then the calculation of the Arnoldi decompositions (11) can be replaced by application of $m$ steps of the symmetric Lanczos process to $\mathbf{A}$ with initial vector $e$. Generically, we obtain
$\mathbf{A}\mathbf{V}_{m}=\mathbf{V}_{m}\mathbf{T}_{m}+{\tilde{v}}_{m+1}e_{m}^{T},$
where the matrix $\mathbf{T}_{m}\in\mathbb{R}^{m\times m}$ is symmetric and tridiagonal, the matrix $\mathbf{V}_{m}\in\mathbb{R}^{n\times m}$ has orthonormal columns with initial column $e/\|e\|_{2}$, and the vector ${\tilde{v}}_{m+1}\in\mathbb{R}^{n}$ is orthogonal to the range of $\mathbf{V}_{m}$; see, e.g., Saad [25] for details on the symmetric Lanczos process.
@@
$L_{f,{\rm Lanczos}}(\mathbf{A},ee^{T})=\mathbf{V}_{m}\mathbf{X}_{m}\mathbf{V}_{m}^{T},$ (17)
where $\mathbf{X}_{m}$ is the upper right $m\times m$ submatrix of the $2m\times 2m$ matrix
$f\left(\left[\begin{array}[]{cc}\mathbf{T}_{m}&ne_{1}e_{1}^{T}\\ \mathbf{0}&\mathbf{T}_{m}\end{array}\right]\right)=\left[\begin{array}[]{cc}f(\mathbf{T}_{m})&\mathbf{X}_{m}\\ \mathbf{0}&f(\mathbf{T}_{m})\end{array}\right];$ (18)
see [16] for further details.
It remains to discuss how to determine approximations of the elements of $L_{f}(\mathbf{A}^{T},ee^{T})$ of largest and smallest magnitude by using the right-hand sides of (12) or (16). We first consider the former. To determine an approximation of an entry of largest magnitude of $L_{f}(\mathbf{A}^{T},ee^{T})$, we first locate an entry $x_{ij}$ of the matrix $\mathbf{X}_{m}$ of largest magnitude and then determine entries of largest magnitude of columns $i$ and $j$ of the matrices $\mathbf{V}_{m}$ and $\mathbf{W}_{m}$, respectively. The product of these entries furnishes an approximation of an entry of $L_{f}(\mathbf{A}^{T},ee^{T})$ of largest magnitude. We proceed analogously to determine an approximation of an entry of $L_{f}(\mathbf{A}^{T},ee^{T})$ of smallest magnitude. Other entries of closest to largest or smallest magnitudes can be computed similarly.
We turn to the use of the right-hand side of (16). To determine an approximation of an entry of largest magnitude of $L_{f}(\mathbf{A}^{T},ee^{T})$, we first locate an entry of the matrix $\frac{f(\mathbf{H}_{m}+hne_{1}e_{1}^{T})-f(\mathbf{H}_{m}-hne_{1}e_{1}^{T})}{2h}$ of largest magnitude. Assume it is entry $\{i,j\}$. Then determine entries of largest magnitude of columns $i$ and $j$ of the matrix $\mathbf{V}_{m}$. The product of these entries yields an approximation of an entry of largest magnitude of $L_{f}(\mathbf{A}^{T},ee^{T})$.
@@ 3 Network modifications based on Perron root sensitivity
The methods of this section require right and left Perron vectors of the adjacency matrix $\mathbf{A}$. When the matrix $\mathbf{A}$ is of small to moderate size, these vectors can be determined with the MATLAB function ${\sf eig}$, which computes all eigenvalues and eigenvectors of $\mathbf{A}$. For large networks, we can compute the Perron vectors with the MATLAB function eigs or with the two-sided Arnoldi method. The latter method was introduced by Ruhe [23] and improved by Zwaan and Hochstenbach [31].
@@ 3.1 Perron communicability for small to medium-sized networks
They are referred to as Perron vectors. Let $\mathbf{F}\in\mathbb{R}^{n\times n}$ be a nonnegative matrix of unit spectral norm, $\|\mathbf{F}\|_{2}=1$, and introduce the small positive parameter $\varepsilon$ and denote the Perron root of $\mathbf{A}+\varepsilon\mathbf{F}$ by $\rho+\delta\rho$. Then
@@
where $\theta$ is the angle between $x$ and $y$. The quantity $1/\cos\theta$ is referred to as the condition number of $\rho$ and denoted by $\kappa(\rho)$; see [29, Section 2]. Note that when $\mathbf{A}$ is symmetric, we have $x=y$, hence $\theta=0$. Equality in (19) is attained when $\mathbf{F}$ is the Wilkinson perturbation $\mathbf{W}=yx^{T}$ associated with $\rho$; see [18, 29] for details.
The total communicability (2) of the graph ${\mathcal{G}}$ can be approximated by the Perron communicability of ${\mathcal{G}}$ [4]:
@@
Typically, $\exp(\rho)$ is a fairly accurate indicator of the Perron communicability and, consequently, of the total communicability. In fact, one has [4]:
@@
Perturb the entry $w_{ij}$ with $i\neq j$ of $\mathbf{A}$ and let

### v2

@@ 2.2 Methods for large networks
Recently, Kandolf et al. KKRS derived a method for evaluating approximations of the Fréchet derivative of a matrix function by Krylov subspace methods. Applications of this technique to network analysis have recently been discussed by De la Cruz Cabrera et al. DLCJNR and Schweitzer Sc . We first outline this method and subsequently discuss some alternatives.
@@
where $\Gamma$ is a curve in $\Omega$ that winds around the spectrum of $\mathbf{A}$ exactly once and $\mathrm{i}=\sqrt{-1}$. In this paper, we are primarily interested in the situation when $f(z)=\exp_{0}(z)$, but the techniques discussed apply to other analytic functions as well. Let $u,v\in\mathbb{R}^{n}$ be nonvanishing vectors. The Fréchet derivative of $f$ at $\mathbf{A}$ in the direction $uv^{T}$ can be expressed as
@@
see, e.g., Hi . Kandolf et al. KKRS determine an approximation of this expression by using Krylov subspace techniques to approximate the vectors
@@
where the superscript H denotes transposition and complex conjugation. Specifically, Kandolf et al. KKRS and Schweitzer Sc approximate the vectors $s(z)$ and $t(z)$ by a Krylov subspace technique based on the Arnoldi process. Application of $1\leq\ell\ll n$ steps of the Arnoldi process to $\mathbf{A}$ with initial vector $u$, and to $\mathbf{A}^{T}$ with initial vector $v$, generically, yields the Arnoldi decompositions
$\mathbf{A}\mathbf{V}_{\ell}=\mathbf{V}_{\ell}\mathbf{G}_{\ell}+{\tilde{v}}_{{\ell}+1}e_{\ell}^{T},\qquad\mathbf{A}^{T}\mathbf{W}_{\ell}=\mathbf{W}_{\ell}\mathbf{M}_{\ell}+{\tilde{w}}_{{\ell}+1}e_{\ell}^{T},$ (12)
@@
where the matrices $\mathbf{G}_{\ell},\mathbf{M}_{\ell}\in\mathbb{R}^{{\ell}\times{\ell}}$ are of upper Hessenberg form, the matrix $\mathbf{V}_{\ell}\in\mathbb{R}^{n\times{\ell}}$ has orthonormal columns with initial column $u/\|u\|_{2}$, the vector ${\tilde{v}}_{{\ell}+1}\in\mathbb{R}^{n}$ is orthogonal to the range of $\mathbf{V}_{\ell}$, the matrix $\mathbf{W}_{\ell}\in\mathbb{R}^{n\times{\ell}}$ has orthonormal columns with initial column $v/\|v\|_{2}$, and the vector ${\tilde{w}}_{{\ell}+1}\in\mathbb{R}^{n}$ is orthogonal to the range of $\mathbf{W}_{\ell}$; see, e.g., Saad (Sa, , Chapter 6) for details on the Arnoldi process. Here we only note that the computation of the decompositions (12) requires the evaluation of ${\ell}$ matrix-vector products with the matrix $\mathbf{A}$ and ${\ell}$ matrix-vector products with the matrix $\mathbf{A}^{T}$. This is the dominating computational work when the matrix $\mathbf{A}$ is large and the number of Arnoldi steps is fairly small. We assume here that the Arnoldi processes do not break down when computing (12); in case of breakdown, the formulas (12) simplify.
Kandolf et al. KKRS propose to use the Arnoldi approximation
$L_{f,{\rm Arnoldi}}(\mathbf{A},uv^{T})=\mathbf{V}_{\ell}\mathbf{X}_{\ell}\mathbf{W}_{\ell}^{T},$ (13)
of $L_{f}(\mathbf{A},uv^{T})$, where $\mathbf{X}_{\ell}$ is the upper right ${\ell}\times{\ell}$ submatrix of the $2{\ell}\times 2{\ell}$ matrix
$f\left(\left[\begin{array}[]{cc}\mathbf{G}_{\ell}&\;\|u\|_{2}\|v\|_{2}\,e_{1}e_{1}^{T}\\ \mathbf{0}&\mathbf{M}_{\ell}^{T}\end{array}\right]\right)=\left[\begin{array}[]{cc}f(\mathbf{G}_{\ell})&\mathbf{X}_{\ell}\\ \mathbf{0}&f(\mathbf{M}_{\ell}^{T})\end{array}\right]$ (14)
@@
with $e_{1}=[1,0,\ldots,0]^{T}$. Schweitzer Sc applies formulas (12), (13), and (14) with $\mathbf{A}$ replaced by $\mathbf{A}^{T}$ and $u=v=e$; cf. (6). Convergence results are provided by Kandolf et al. KKRS .
@@
An alternative approach to compute an approximation of the Fréchet derivative $L_{f}(\mathbf{A}^{T},ee^{T})$ is to apply finite-difference approximations analogously as (10). Application of ${\ell}$ steps of the Arnoldi process to the matrix $\mathbf{A}^{T}$ with initial vector $e$ gives the Arnoldi decomposition
$\mathbf{A}^{T}\mathbf{U}_{\ell}=\mathbf{U}_{\ell}\mathbf{H}_{\ell}+{\tilde{u}}_{{\ell}+1}e_{\ell}^{T},$
@@
where the columns of $\mathbf{U}_{\ell}\in\mathbb{R}^{n\times{\ell}}$ are orthonormal and span the Krylov subspace
$\mathbf{K}_{{\ell}}(\mathbf{A}^{T},e)={\rm span}\{e,\mathbf{A}^{T}e,(\mathbf{A}^{T})^{2}e,\ldots,(\mathbf{A}^{T})^{{\ell}-1}e\}$
@@
and the first column of $\mathbf{U}_{\ell}$ is $e/\sqrt{n}$. Moreover, the vector ${\tilde{u}}_{{\ell}+1}\in\mathbb{R}^{n}$ is orthogonal to $\mathbf{K}_{\ell}(\mathbf{A}^{T},e)$ and $\mathbf{H}_{\ell}\in\mathbb{R}^{{\ell}\times{\ell}}$ is an upper Hessenberg matrix. Then
$\mathbf{U}_{\ell}^{T}\mathbf{A}^{T}\mathbf{U}_{\ell}=\mathbf{H}_{\ell}.$
@@
$\mathbf{A}^{T}\approx\mathbf{U}_{\ell}\mathbf{H}_{\ell}\mathbf{U}_{\ell}^{T}$ (15)
@@
$f(\mathbf{A}^{T})\approx\mathbf{U}_{\ell}f(\mathbf{H}_{\ell})\mathbf{U}_{\ell}^{T}.$ (16)
The approximation (16) is quite accurate when $f(\mathbf{A}^{T})$ can be approximated well by a matrix of low rank. This is the case for many real-life undirected networks when $f(t)=\exp(t)$; see FMRR .
@@
$\mathbf{A}^{T}+hee^{T}\approx\mathbf{U}_{\ell}\mathbf{H}_{\ell}\mathbf{U}_{\ell}^{T}+hee^{T}=\mathbf{U}_{\ell}(\mathbf{H}_{\ell}+hne_{1}e_{1}^{T})\mathbf{U}_{\ell}^{T}$
@@
$f(\mathbf{A}^{T}+hee^{T})\approx\mathbf{U}_{\ell}f(\mathbf{H}_{\ell}+hne_{1}e_{1}^{T})\mathbf{U}_{\ell}^{T},$
@@
$\frac{f(\mathbf{A}^{T}+hee^{T})-f(\mathbf{A}^{T}-hee^{T})}{2h}\approx\mathbf{U}_{\ell}\frac{f(\mathbf{H}_{\ell}+hne_{1}e_{1}^{T})-f(\mathbf{H}_{\ell}-hne_{1}e_{1}^{T})}{2h}\mathbf{U}_{\ell}^{T}.$ (17)
We will use the expression on the right-hand side as an approximation of $L_{f}(\mathbf{A}^{T},ee^{T})$ in computed examples with $h$ the same as in Example 2.1. Note that the evaluation of this expression only requires the computation of $\ell$ matrix-vector products with the matrix $\mathbf{A}^{T}$. For large-scale problems for which the evaluation of matrix-vector products is the dominant computational work, under the assumption that both methods require the same number of iterations, the use of the right-hand side of (17) halves the computational burden when compared with the evaluation of (13).
We turn to the situation when the matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$ is symmetric and first review the computations described by Kandolf et al. KKRS of the analogue of the expression (13) when the direction is $ee^{T}$. Then the calculation of the Arnoldi decompositions (12) can be replaced by application of ${\ell}$ steps of the symmetric Lanczos process to $\mathbf{A}$ with initial vector $e$. Generically, we obtain
$\mathbf{A}\mathbf{U}_{\ell}=\mathbf{U}_{\ell}\mathbf{T}_{\ell}+{\tilde{u}}_{{\ell}+1}e_{\ell}^{T},$
where the matrix $\mathbf{T}_{\ell}\in\mathbb{R}^{{\ell}\times{\ell}}$ is symmetric and tridiagonal, the matrix $\mathbf{U}_{\ell}\in\mathbb{R}^{n\times{\ell}}$ has orthonormal columns with initial column $e/\|e\|_{2}$, and the vector ${\tilde{v}}_{{\ell}+1}\in\mathbb{R}^{n}$ is orthogonal to the range of $\mathbf{U}_{\ell}$; see, e.g., Saad Sa0 for details on the symmetric Lanczos process.
@@
$L_{f,{\rm Lanczos}}(\mathbf{A},ee^{T})=\mathbf{U}_{\ell}\mathbf{X}_{\ell}\mathbf{U}_{\ell}^{T},$ (18)
where $\mathbf{X}_{\ell}$ is the upper right ${\ell}\times{\ell}$ submatrix of the $2{\ell}\times 2{\ell}$ matrix
$f\left(\left[\begin{array}[]{cc}\mathbf{T}_{\ell}&ne_{1}e_{1}^{T}\\ \mathbf{0}&\mathbf{T}_{\ell}\end{array}\right]\right)=\left[\begin{array}[]{cc}f(\mathbf{T}_{\ell})&\mathbf{X}_{\ell}\\ \mathbf{0}&f(\mathbf{T}_{\ell})\end{array}\right];$ (19)
see KKRS for further details.
It remains to discuss how to determine approximations of the elements of $L_{f}(\mathbf{A}^{T},ee^{T})$ of largest and smallest magnitude by using the right-hand sides of (13) or (17). We first consider the former. To determine an approximation of an entry of largest magnitude of $L_{f}(\mathbf{A}^{T},ee^{T})$, we first locate an entry $x_{ij}$ of the matrix $\mathbf{X}_{\ell}$ of largest magnitude and then determine entries of largest magnitude of columns $i$ and $j$ of the matrices $\mathbf{V}_{\ell}$ and $\mathbf{W}_{\ell}$, respectively. The product of these entries furnishes an approximation of an entry of $L_{f}(\mathbf{A}^{T},ee^{T})$ of largest magnitude. We proceed analogously to determine an approximation of an entry of $L_{f}(\mathbf{A}^{T},ee^{T})$ of smallest magnitude. Other entries of closest to largest or smallest magnitudes can be computed similarly.
We turn to the use of the right-hand side of (17). To determine an approximation of an entry of largest magnitude of $L_{f}(\mathbf{A}^{T},ee^{T})$, we first locate an entry of the matrix $\frac{f(\mathbf{H}_{\ell}+hne_{1}e_{1}^{T})-f(\mathbf{H}_{\ell}-hne_{1}e_{1}^{T})}{2h}$ of largest magnitude. Assume it is entry $\{i,j\}$. Then determine entries of largest magnitude of columns $i$ and $j$ of the matrix $\mathbf{U}_{\ell}$. The product of these entries yields an approximation of an entry of largest magnitude of $L_{f}(\mathbf{A}^{T},ee^{T})$.
@@ 3 Network modifications based on Perron root sensitivity
The methods of this section require right and left Perron vectors of the adjacency matrix $\mathbf{A}$. When the matrix $\mathbf{A}$ is of small to moderate size, these vectors can be determined with the MATLAB function ${\sf eig}$, which computes all eigenvalues and eigenvectors of $\mathbf{A}$. For large networks, we can compute the Perron vectors with the MATLAB function eigs or with the two-sided Arnoldi method. The latter method was introduced by Ruhe Ru and improved by Zwaan and Hochstenbach ZH .
@@ 3.1 Perron communicability for small to medium-sized networks
They are referred to as Perron vectors. Let $\mathbf{F}\in\mathbb{R}^{n\times n}$ be a nonnegative matrix of unit spectral norm, $\|\mathbf{F}\|_{2}=1$. Introduce the small positive parameter $\varepsilon$ and denote the Perron root of $\mathbf{A}+\varepsilon\mathbf{F}$ by $\rho+\delta\rho$. Then
@@
where $\theta$ is the angle between $x$ and $y$. The quantity $1/\cos\theta$ is referred to as the condition number of $\rho$ and denoted by $\kappa(\rho)$; see (Wi, , Section 2). Note that when $\mathbf{A}$ is symmetric, we have $x=y$, hence $\theta=0$. Equality in (20) is attained when $\mathbf{F}$ is the Wilkinson perturbation $\mathbf{W}=yx^{T}$ associated with $\rho$; see MSN ; Wi for details.
The total communicability (2) of the graph ${\mathcal{G}}$ can be approximated by the Perron communicability of ${\mathcal{G}}$ DLCJNR :
@@
Typically, $\exp(\rho)$ is a fairly accurate indicator of the Perron communicability and, consequently, of the total communicability. In fact, one has DLCJNR :
@@
Perturb the entry $w_{ij}$ with $i\neq j$ of $\mathbf{A}$ by $\varepsilon\neq 0$ and let